OpenStudy (kkutie7):

I need help finding the general term of a taylor series assuming value of the index is k=0 cos(t)=-1+((t-pi)^2)/2-((t-pi)^4)/24+((t-pi)^6)/720+....

OpenStudy (kkutie7):

\[\sum_{k=0}^{\infty}\frac{ (-1)^{k+1} (t-\pi)^{2k}}{ ? }\]

OpenStudy (dan815):

(2k)!

OpenStudy (kkutie7):

dang that was way easier than I thought

OpenStudy (dan815):

=]